OpenStudy (anonymous):

-x + 3 = y for y = mx + 6

OpenStudy (anonymous):

m = -2

OpenStudy (anonymous):

Your answer for x will probably be in terms of m

OpenStudy (anonymous):

cowsgomoo, I know what you did. You multiplied 2 into the equation -x+3=y but it's not correct cause it'll make y into 2y. Which means the simultaneous equation doesn't work.

OpenStudy (anonymous):

if you need to find m, then \[m = (-x-3)/x\]

OpenStudy (anonymous):

if you need to find x, then \[x = 3/(-1-m)\]

OpenStudy (anonymous):

turabq, if you multiply the equation by 2, the equation will be -2x+6=2y. You can't solve that simultaneously with y=mx+6

OpenStudy (anonymous):

@Japorized right, i need a break.